# Group-Adaptive Threshold Optimization for Robust AI-Generated Text Detection

### v3

@@ Abstract
The advancement of large language models (LLMs) has made it difficult to differentiate human-written text from AI-generated text. Several AI-text detectors have been developed in response, which typically utilize a fixed global threshold (e.g., $\theta=0.5$) to classify machine-generated text. However, we find that one universal threshold can fail to account for subgroup-specific distributional variations. For example, when using a fixed threshold, detectors make more false positive errors on shorter human-written text than longer, and more positive classifications on neurotic writing styles than open among long text. These discrepancies can lead to misclassification that disproportionately affects certain groups. We address this critical limitation by introducing FairOPT, an algorithm for group-specific threshold optimization in AI-generated content classifiers. Our approach partitions data into subgroups based on attributes (e.g., text length and writing style) and learns decision thresholds for each group, which enables careful balancing of performance and fairness metrics within each subgroup. In experiments with four AI text classifiers on three datasets, FairOPT enhances overall F1 score and decreases balanced error rate (BER) discrepancy across subgroups. Our framework paves the way for more robust and fair classification criteria in AI-generated output detection. We release our data, code, and project information at URL.
detection, robustness, fairness, threshold optimization
[…]
mO May[#1]
@@ 1 Introduction
The large-scale adoption of large language models (LLMs) has led to the widespread dissemination of AI-generated text that is challenging to differentiate from human-written content (Wu et al., 2025). This has raised concerns around the spread of misinformation (Chen & Shu, 2024), the degradation of publication standards (Wu et al., 2023), potential cybersecurity threats (Yao et al., 2024), and breaches of academic integrity (Perkins, 2023). As a result, effective AI content detection systems are essential for mitigating the risks from generative models. Recently, there has been substantial development in AI text detection tools, such as RoBERTa-based models (Solaiman et al., 2019), DetectGPT (Mitchell et al., 2023), Fast-DetectGPT (Bao et al., 2023), GLTR (Gehrmann et al., 2019), RADAR (Hu et al., 2023), and efforts to use LLMs as detectors (Bhattacharjee & Liu, 2024). Generally, these methods estimate a probability $p$ that the text is generated by AI by comparing the patterns (e.g., token probabilities or logits) of the given text against AI-generated text, and use a single universal threshold of $p>0.5$ for classification (Freeman & Moisen, 2008).
@@
Figure 1: We find that using one universal probability threshold (often $\theta>0.5$) to classify AI and human-generated text can fail to account for subgroup-specific distributional variations, such as text length or writing style, leading to higher errors for certain demographics. We propose FairOPT to learn multiple decision thresholds tailored to each subgroup, which leads to more robust AI-text detection.
@@
Figure 2: Applying adaptive thresholds to different probability distributions by subgroup, as seen through cumulative density functions (CDFs). We partition texts by three length categories (short, medium, long) and five personality traits (extroversion, neuroticism, agreeableness, conscientiousness, openness), and use RADAR (Hu et al., 2023) to infer AI-generated probabilities on the test dataset. Observe that a static classification threshold ($0.5$, in black) and a single optimized threshold (in gray at the right end side) does not account for subgroup-specific distributional variations. We used the thresholds in Appendix C. The static threshold followed conventional $0.5$, AUROC method is noted in Appendix B.
However, a single universal threshold for AI content classification fails to account for text with different characteristics, such as length or writing style, which can have very different probability distributions. That is to say, one fixed threshold does not consider the variation of the probability. For example, when using a fixed threshold of 0.5, a writing style characterized by openness and positivity is more likely to be identified as AI-generated than one characterized by neuroticism, as illustrated in Figure 2. This can lead to misclassification and disproportionately affect certain groups. Similar concerns were raised regarding race (Alghamdi et al., 2022; Bendekgey & Sudderth, 2021) and gender (Weber et al., 2020; Jang et al., 2022; Bendekgey & Sudderth, 2021).
To address these limitations, we propose FairOPT to learn decision thresholds for distinct subgroups rather than using a standard threshold (see Figure 1). Our method balances performance (e.g., ACC, F1, precision) and fairness metrics (e.g., demographic parity, equality of odds, equal opportunity) in learning these thresholds, ensuring that predefined fairness criteria are met across all subgroups in question. In experiments with four AI text classifiers on three datasets, FairOPT increases overall F1 score and decreases balanced error rate (BER) discrepancy across subgroups, yielding better overall tradeoffs between performance and fairness. To summarize, our main contributions are: (a) identifying discrepancies in AI-generated text classifiers based on length and stylistic characteristics, (b) developing FairOPT, a group-adaptive decision threshold method to enhance group performance and fairness, and (c) developing a model with a rule-based relaxed fairness criteria to achieve fast convergence even when perfect fairness is not possible.
@@ 2 Related Work
We review work in AI-generated text detection and methods to threshold probability distributions into classifier labels.
AI text detectors
AI text detectors have been developed to flag AI-generated text. RADAR utilizes adversarial learning techniques (Hu et al., 2023), GLTR employs statistical metrics, including entropy, distributions, and ranks, to discern anomalies suggestive of AI content (Gehrmann et al., 2019), DetectGPT examines the curvature of log probability distributions, concentrating on local likelihood curvature to detect AI-generated text (Mitchell et al., 2023), and RoBERTa-based detectors fine-tune pretrained language models to proficiently categorize text as either human-produced or AI-generated (Solaiman et al., 2019). We employed probabilistic classifiers among them which require a threshold-based decision mechanism. Unlike direct discrimination methods, thresholding enables post-processing for the balancing of performance and fairness.
Fixed universal thresholds for classification:
@@
The most prevalent approach for making decision thresholds in classification involves employing a fixed universal threshold of $0.5$ to map predicted probabilities to class labels (Freeman & Moisen, 2008). It is commonly used in many fields like photogrammetry (Shao et al., 2016), ecology (Manel et al., 1999; Hanberry & He, 2013) and computer science (Lu et al., 2024). Nevertheless, as highlighted by Freeman & Moisen (2008), the dependence on $\theta=0.5$ as a default threshold, merely due to its general acceptance, is frequently unreliable owing to variations in data distributions.
Universal thresholds with optimization:
@@
Threshold optimization techniques have been formulated to learn optimal thresholds across entire datasets. These methods optimize metrics such as the area under the receiver operating characteristic curve (AUROC) (Bradley, 1997) and derive optimized decision thresholds under constraints such as the false positive rate (FPR) (Krishna et al., 2024; Lipton et al., 2014). These approaches find applications across disciplines, including econometrics (Staňková, 2023), statistics (Esposito et al., 2021), and machine learning (OpenAI, 2023). While these methods yield a globally optimized threshold tailored to the chosen model, they can fail to adapt to the characteristics of individual instances, which can compromise the robustness of detection systems, particularly in significantly divergent data distributions.
Adaptive thresholds across groups:
@@
Recent work has highlighted the necessity for adaptive thresholds across groups (Jang et al., 2022; Bakker et al., 2021). Menon & Williamson (2018) propose instance-dependent thresholds for inferred probabilities, resulting in optimal classifiers under cost-sensitive fairness objectives and enhancing overall performance. Similarly, Corbett-Davies et al. (2017) demonstrate that group-specific thresholds, informed by group-level statistics, can facilitate fairness-aware classification. Canetti et al. (2019) investigate thresholds differentiated by race and proved that adaptive thresholds reduce performance disparity. Jang et al. (2022) present threshold-adaptation methods to ensure fair classification, and Bakker et al. (2021) explore threshold tuning to maintain stable classification performance across subgroups. Our method, FairOPT is grounded in the aforementioned principles and introduce a novel approach of adaptive thresholding.
@@ 3 Conceptual Framework and Methodology
This section outlines our method for adaptive thresholds in AI-generated text detection, FairOPT, which iteratively adjusts classification thresholds for different subgroups to balance performance with inter-group fairness.
@@ 3.1 Objective
Considering a set of subgroups $G_{1},\ldots,G_{i}$ characterized by attributes (e.g., length and style) and acknowledging that texts with varying characteristics shows discrepancies in probability distributions, our objective is to refine the binary classification thresholds $\{\theta(G_{1}),\ldots,\theta(G_{i})\}$ applicable to $G_{1},\ldots,G_{i}$. By accounting for subgroup-specific variations, these tailored thresholds ensure that decision boundaries accurately reflect the specific probability distributions.
@@ 3.1.1 Preliminaries of AI Text Detection
The detection of AI-generated text is commonly formulated as a binary classification task, where the input space $\mathcal{X}$ consists of the sequence of tokenized indices derived from a given text, and the label space $\mathcal{Y}=\{0,1\}$ denotes the two distinct classes of 00 for human-written text and $1$ for AI-generated text. To address this task, existing work primarily utilizes a pretrained neural network as a probabilistic classifier $M_{\theta}:\mathcal{X}\rightarrow[0,1]$, which assigns a probability $p$ to the input reflecting the likelihood of it being AI-generated (Solaiman et al., 2019; Hu et al., 2023). This probability is then converted into a predicted label by applying a threshold, typically set at $0.5$. Thus, the final predicted label $\hat{Y}\in\{0,1\}$ is determined as follows:
@@ 3.1.2 Performance metrics
After inferring $\hat{Y}$, we can compare $Y$ and $\hat{Y}$. Correctly classifying AI-generated content as AI constitutes a true positive (TP). If human text is not classified as AI, it is true negative (TN). Flagging human text as AI is a false positive (FP), and flagging AI text as human is a false negative (FN). Details about the classification matrix and performance metrics are in Appendix A.1.
@@ 3.1.3 Fairness metrics
The objective of fair classification is to mitigate disparities while preserving performance. To quantify fair performance, we focus on Demographic Parity (DP) (Kim et al., 2020), a fairness criterion measuring whether a classifier’s predictions $\hat{Y}$ are statistically independent across attributes $S$: $\hat{Y}\perp\!\!\!\perp S$. This condition mandates that the probability of assigning a positive outcome is equitable across all groups, regardless of the true class labels $Y$. To measure the extent of disparity, we define the DP metric:
@@
By addressing an overall equal rate of positive predictions across groups, DP seeks to measure disparate impact, where one group receives a disproportionate share of positive or negative outcomes due to biases in the classifier model (Feldman et al., 2015). In applying DP, the perfect fairness is $\Delta_{\text{DP}}=0$, which means that there is no difference in positive rates between groups. Details about fairness and performance metrics are noted in Appendix A.2.
@@ 3.2 Relaxed Fairness
While achieving 100% fairness across groups is theoretically ideal, it is often impractical due to computational constraints and the difficulty of ensuring model convergence under strict fairness criteria (Dwork et al., 2012). The 80% fairness rule, often called relaxed fairness, offers a pragmatic approach to balance reasonable fairness and performance (Feldman et al., 2015), and can also mitigate overfitting risks and ensure convergence based on the rule-based constraints.
We formalize a relaxed fairness criterion requiring that the ratio of minimum to maximum value of each fairness metric across all subgroups does not fall below a specified threshold $\tau$, set to $0.8$. It can be mathematically represented as:
@@
  • $\tau\in[0,1]$ is the fairness criteria, with $\tau=0.8$ reflecting the 80% rule and $\tau=1$ reflects 100% fairness.
Note that we should consider limiting the number of fairness metrics because of the impossibility theorem (Dwork et al., 2012) as it is computationally impractical to satisfy all fairness criteria simultaneously. The generated thresholds for subgroups based on the training dataset is used for the classification criteria in the test dataset that has same subgroups.
3.3 Group-Adaptive Threshold Optimization
Process
Given these preliminaries, we now introduce our FairOPT algorithm, which is summarized in Algorithm 1. At a high level, FairOPT uses a gradient-based subgroup threshold update to jointly optimize classification performance (ACC and F1) and group-level fairness.
Formally, we partition the dataset $D$ into subgroups $\{G_{1},\dots,G_{i}\}$ based on features $\{S_{1},\dots,S_{n}\}$. Each subgroup $G_{i}$ is assigned an initial threshold $\theta_{\mathrm{init}}$ (e.g., 0.5 to all groups).
During each iteration, we compute predictions for each sample $x_{n}\in G_{i}$ by binarizing the predicted probability $p_{n}$ using the current threshold $\theta(G_{i})$. The model derives a contingency matrix based on the predictions $\hat{y}_{n}$ and actual values $y_{n}$.
For each subgroup $G_{i}$, we define a loss function that penalizes both low accuracy and insufficient F1 score:
where $\beta$ is a minimum F1 threshold, $\kappa>0$ is a penalty weight, $\alpha$ is also given for the minimum ACC, and $[\cdot]_{+}$ denotes the positive part that only positive values are retained, turning negative numbers to zero. The minimum thresholds are essential for ensuring stable convergence and preventing degenerate outcomes across all subgroups.
The finite-difference approximation estimates the gradient of each subgroup’s threshold-based loss function by computing the difference between the loss values at $\theta(G_{i})+\delta$ and $\theta(G_{i})-\delta$, then dividing by $2\delta$ (Liu et al., 2020). This method is employed for an effective interaction with the discrete contingency matrix, as it allows gradient estimation without requiring an explicit analytic form of the loss function’s derivative.
We then perform a gradient descent step for each subgroup:
followed by clipping to the allowable interval $\bigl{[}a,b\bigr{]}\subseteq[0,1]$ to make sure it is a valid probability.
@@ 3.3 Group-Adaptive Threshold Optimization
10:  Output: $\{\theta(G_{1}),\dots,\theta(G_{j})\}$.
Finally, the model calculates domain-relevant fairness metrics $\{M_{1},\ldots,M_{k}\}$ across all subgroups by measuring the maximum inter-group disparity:

### v4

@@ Abstract
The advancement of large language models (LLMs) has made it difficult to differentiate human-written text from AI-generated text. Several AI-text detectors have been developed in response, which typically utilize a fixed global threshold (e.g., $\theta=0.5$) to classify machine-generated text. However, one universal threshold could fail to account for distributional variations by subgroups. For example, when using a fixed threshold, detectors make more false positive errors on shorter human-written text, and more positive classifications of neurotic writing styles among long texts. These discrepancies can lead to misclassifications that disproportionately affect certain groups. We address this critical limitation by introducing FairOPT, an algorithm for group-specific threshold optimization for probabilistic AI-text detectors. We partitioned data into subgroups based on attributes (e.g., text length and writing style) and implemented FairOPT to learn decision thresholds for each group to reduce discrepancy. In experiments with nine AI text classifiers on three datasets, FairOPT decreases overall balanced error rate (BER) discrepancy by 12% while minimally sacrificing accuracy by 0.003%. Our framework paves the way for more robust classification in AI-generated content detection via post-processing. We release our data, code, and project information at URL.
@@ 1 Introduction
The copious use of large language models (LLMs) facilitated widespread dissemination of AI-generated text [1, 2] and has raised concerns around the dissemination of misinformation [3], the degradation of publication standards [4], cybersecurity threats [5], and breaches of academic integrity [6]. In response to the anticipated risks, AI detection techniques have been proposed [1, 7]. Typically, these methods estimate a probability $p$ that the text is generated by AI by comparing the patterns (e.g., token probabilities) of the given text against AI-generated text, and use a universal threshold (e.g., $\theta>0.5$[8] for a binary classification.
@@
Figure 1: We find that using one universal probability threshold (often $\theta>0.5$) to classify AI and human-generated text can fail to account for subgroup-specific distributional variations, such as text length or writing style, leading to higher errors for certain demographics. We propose FairOPT to learn multiple decision thresholds tailored to each subgroup, which reduces discrepancy in AI text detection as shown in Figure 7.
@@
Figure 2: Applying adaptive thresholds to different probability distributions by subgroup in the test dataset, as seen through cumulative density functions (CDFs). We partition texts by three length categories and five personality traits as noted in Table 1. Observe that a static classification threshold ($0.5$, in black) and a single optimized threshold using AUROC-based method (in gray) do not account for subgroup-specific distributional variations. We used fifteen thresholds that have been generated using FairOPT to address the varying probabilistic distribution from AI text classifiers. The training is based on the RoBERTa-large detector. Details about the thresholds are in Appendix E.
However, a single universal threshold for AI content classification neglects the heterogeneity of probability distributions by textual characteristics, such as length and writing style. To put in another way, one fixed threshold does not consider the variation of the probability from various stylistic features. For example, when using a fixed threshold, a writing style characterized by long neurotic writing could be more likely to be identified as AI-generated than one characterized by agreeableness, resulting in disproportionate misclassifications, affecting certain groups. Analogous disparities have been documented in other classification tasks, regarding race [9, 10] and gender [11, 12, 10].
To overcome these limitations, we introduce FairOPT, a group-adaptive thresholding framework that learns distinct decision thresholds for subgroups rather than using methods with one universal threshold (see Figure 1). Our method balances performance (e.g., accuracy (ACC), F1, precision) and fairness metrics (e.g., demographic parity, or equality of odds) in learning thresholds. Experiment on nine AI text detectors on three benchmark datasets demonstrates that FairOPT substantially reduces BER disparity across subgroups while incurring only a marginal tradeoff in overall performance.
To summarize, our main contributions are: (a) identifying significant discrepancies in AI-generated text classifiers based on text length and stylistic characteristics, (b) developing FairOPT, a group-adaptive decision threshold that calibrates the decision boundaries to reduce group discrepancy, and (c) showing the effectiveness of the FairOPT, achieving a 12% reduction in discrepancy while showing negligible tradeoff in performance by dropping less than 0.1% of ACC.
@@ 2 Related Work
AI text detectors:
AI text detectors have been developed to tackle issues related to misinformation, integrity, and security. RADAR utilizes adversarial learning techniques [13], GLTR employs statistical metrics, including entropy, probability distribution over the model’s vocabulary for each next token (i.e. the softmax scores across all candidate tokens), and records the rank position of the actual chosen token within that sorted probability to discern [14], DetectGPT examines the curvature of log probability distributions of the referred LLM, concentrating on local likelihood curvature to detect AI-generated text [15], and RoBERTa-based detectors fine-tune pretrained language models [16]. We employed probabilistic classifiers for an AI-likelihood score that use a threshold for classification rather than architectures that generate a direct discrimination.
Static thresholding for classification:
@@
The canonical thresholding in probabilistic classification is using a threshold of $0.5$ to map the probabilities to labels [8]. This default rule is commonly used in many fields like photogrammetry [17], ecology [18, 19], and machine learning [20]. Nevertheless, as highlighted by Freeman and Moisen [8], the dependence solely on $\theta=0.5$ as a default threshold, merely due to its conventional acceptance, could be unreliable because of variations in data.
Universal thresholding with optimization:
@@
Threshold optimization techniques have been formulated to learn an optimal threshold across the dataset. These methods optimize classifiers’ performance metrics, such as the True Positive rate (TPR) and false positive rate (FPR) [21, 22], and the most representative method is optimizing using the receiver operating characteristic curve (AUROC) [23]. These approaches find applications across disciplines, including econometrics [24], statistics [25], and machine learning [26]. While the AUROC-based method yields a globally optimized threshold tailored to the chosen model, it can lead to over-fitting or fail to adapt to other models.
Adaptive thresholding across groups:
@@
Several studies advocate for group-adaptive decision thresholds to mitigate disparities [12, 27]. Menon and Williamson [28] propose instance-dependent thresholds for inferred probabilities, resulting in reduced discrepancy. Similarly, Corbett-Davies et al. [29] demonstrate that group-specific thresholds can facilitate fairness-aware classification. Canetti et al. [30] investigate thresholds differentiated by race and proved that adaptive thresholds reduce performance disparity. Jang et al. [12] present threshold-adaptation methods to ensure fair classification, and Bakker et al. [27] explore threshold tuning to maintain stable classification performance. Building on these foundations, our framework FairOPT introduces a novel group-adaptive thresholding algorithm that jointly optimizes performance and fairness objectives.
Figure 3: Human-written short texts have a higher FPR compared to other text lengths, given the static threshold, showing the inadequacy of the static threshold to variations. Curves reflect the average probability of AI detectors in the training dataset.
@@ 3 Conceptual Framework and Methodology
This section outlines our method for adaptive thresholding, FairOPT, which iteratively adjusts classification thresholds for different subgroups to reduce discrepancy.
@@ 3.1 Objective
Consider a set of subgroups $G_{1},\ldots,G_{i}$ with attributes (e.g., length and style). Acknowledging that texts with varying characteristics show variations in probability distributions, our objective is to find optimal thresholds $\{\theta(G_{1}),\ldots,\theta(G_{i})\}$ for binary classification.
@@ 3.1.1 Preliminaries of AI Text Detection
The detection of AI-generated text is commonly formulated as a probabilistic classification for a binary classification, where the input space $\mathcal{X}$ consists of the sequence of tokens derived from a given text, and the label space $\mathcal{Y}=\{0,1\}$ denotes the two distinct classes of 00 for human-written and $1$ for AI-generated. To address this task, existing work primarily utilizes a pretrained neural network as a probabilistic classifier $M_{\theta}:\mathcal{X}\rightarrow[0,1]$, which assigns a probability $p$ to the input reflecting the likelihood of it being AI-generated [16, 13]. This probability is then converted into a predicted label by applying a threshold, typically set at $0.5$. Thus, the predicted label, AI-written, $\hat{Y}\in\{0,1\}$ is determined as follows:
@@ 3.1.2 Performance metrics
After inferring $\hat{Y}$, we can compare $Y$ and $\hat{Y}$. Details about the confusion matrix and performance metrics are in Appendix B.1.
@@ 3.1.3 Fairness metrics
To quantify disparity, we use Demographic Parity (DP) [31] and Balanced Error Rate (BER) [32]. DP is a fairness criterion measuring whether a classifier’s predictions $\hat{Y}$ are statistically independent across attributes $S$: $\hat{Y}\perp\!\!\!\perp S$. This condition mandates that the probability of assigning a positive outcome is equal across all groups. We define the DP metric:
@@
DP seeks to measure disparate impact, where one group receives disproportionate results [33], by addressing an overall equal rate of positive predictions, encompassing TP and TN, across groups. In applying DP, the perfect fairness is $\Delta_{\text{DP}}=0$, which means no difference in overall TP and TN between groups.
In our subgroup-adaptive evaluation, we use BER to equally measure the type I and II errors, which are FP and FN, while including positive predictions from DP. For each subgroup $G_{i}$ we compute its BERi to measure the discrepancy, and the lower discrepancy indicates more equitable classification across groups.
$\mathrm{BER}=\tfrac{1}{2}\Bigl{(}\underbrace{\tfrac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TN}}}_{\mathrm{FPR}}\;+\;\underbrace{\tfrac{\mathrm{FN}}{\mathrm{FN}+\mathrm{TP}}}_{\mathrm{FNR}}\Bigr{)}$ (3)
$\text{Discrepancy}=\max\mathrm{BER}_{i}\;-\;\min\mathrm{BER}_{i}$ (4)
Details about fairness metrics are noted in Appendix B.2.
Figure 4: The FairOPT algorithm enables us to obtain a balance in the tradeoff between performance and fairness, through a relaxed fairness criterion discussed in Section 3.2. Details about the tradeoff are in Appendix H.2.
@@ 3.2 Relaxed Fairness
While achieving $100\%$ fairness across groups (e.g., $DP=0$ across fifteen groups) is theoretically ideal, it is often impractical due to computational constraints and the difficulty of ensuring model convergence under strict fairness criteria [34]. The $80\%$, relaxed fairness, rule has been suggested to offer a pragmatic approach (e.g., allowing $DP=0.2$ in achieving fairness) [33], which offers a more feasible fairness criterion and mitigates overfitting risks.
We formalize a relaxed fairness criterion requiring that the ratio of minimum to maximum value of each fairness metric across all subgroups does not fall below a given flexibility $\tau$, set to $0.8$:
@@
  • $\tau\in[0,1]$ is the fairness criteria, with $\tau=1$ reflects 100% fairness, which is the most strict.
We limited the number of fairness metrics due to impossibility theorem [34] to make it computationally feasible, as it is impractical to satisfy all fairness criteria simultaneously.
Algorithm 1 FairOPT: Group-Specific Thresholding
@@
10:        end if
11:     end for
12:  end for
13:  Output: $\{\theta(G_{1}),\dots,\theta(G_{j})\}$.
3.3 Group-Adaptive Threshold Optimization
Process
Given this setup, we introduce FairOPT  which is described in Algorithm 1. FairOPT uses a gradient-based subgroup threshold update to jointly optimize group-level fairness and classification performance. Formally, we partition the dataset $D$ into subgroups $\{G_{1},\dots,G_{i}\}$ based on features $\{S_{1},\dots,S_{n}\}$. Each subgroup $G_{i}$ is assigned an initial threshold $\theta_{\mathrm{init}}$ (e.g., 0.5 to all groups). During each iteration, we compute predictions for each sample $x_{n}\in G_{i}$ by binarizing the predicted probability $p_{n}$ using the current threshold $\theta(G_{i})$. The model derives a contingency matrix based on the predictions $\hat{y}_{n}$ and actual values $y_{n}$. For each subgroup $G_{i}$, we define a loss function that penalizes both low accuracy and insufficient F1 score:
where $\beta$ is a minimum F1, $\kappa>0$ is a penalty weight, $\alpha$ is given for the minimum ACC, and $[\cdot]_{+}$ denotes the positive part that only positive values are retained, turning negative numbers to zero. These minimum standards are essential for ensuring stable convergence and preventing degenerate outcomes across all subgroups.
The finite-difference approximation estimates the gradient of each subgroup’s threshold-based loss function by computing the difference between the two loss values at $\theta(G_{i})+\delta$ and $\theta(G_{i})-\delta$, then dividing by $2\delta$ [35].
We then perform a gradient descent step for each subgroup:
followed by clipping to $\bigl{[}a,b\bigr{]}\subseteq[0,1]$ to ensure a valid probability.
Finally, the model calculates the given fairness metric(s) $\{M_{1},\ldots,M_{k}\}$ across all subgroups by measuring the inter-group disparity: